The length of a rectangle is 3 more than twice its width. The perimeter is 48 feet. Find the width.
Answer by lazi(15)

L = 2W + 3
P = 48
P = 2(W + L)
48 = 2(w + 2W + 3)
48 = 2(3W + 3)
48 = 2 * 3(W + 1)
48 = 6(W + 1)
8 = W + 1
7 = W